# Fibre homogenisation for time-dependent problems

### v1

@@ 1 Introduction
This article is concerned with the quantitative homogenisation of evolutionary equations with rapidly oscillating periodic spatial coefficients
@@
where $M$ and $N$ are bounded and positivity coefficient matrices, the small parameter $\varepsilon$ is the spatial period and $A$ is some first order differential operator. The usual, now classical, ‘qualitative’ homogenisation theory states that the solution $U_{\varepsilon}$ converges (in some function space), as $\varepsilon$ tends to zero, to $U_{0}$ the solution to a evolutionary equations with constant coefficients
@@
for $U_{0}$ the solution to the homogenised evolutionary wave system
@@
$\partial_{t}u_{0}-\mathrm{div}(b_{0}\nabla u_{0})=f,\quad t\in\mathbb{R},\,x\in\mathbb{R}^{d},$
where $b_{0}$ is the homogenised matrix (associated to $a$). As in Section 2, we additionally have the corrector estimate
@@
For a given Hilbert space $H$ with inner product $\langle\cdot,\cdot\rangle_{H}$ and norm $\|\cdot\|_{H}:=\langle\cdot,\cdot\rangle^{1/2}$, we denote: $L(H)$ to be the space of bounded linear operators in $H$,
@@
and $H_{\nu}^{1}(\mathbb{R};H)$ is the space of $H$-valued, weakly differentiable $L_{2,\nu}(\mathbb{R};H)$-functions with distributional derivative representable as an $L^{2}_{\nu}(\mathbb{R};H)$-function. $L^{2}_{\nu}(\mathbb{R};H)$ and $H^{1}_{\nu}(\mathbb{R};H)$ are Hilbert spaces with the standard inner products. For a given function space $X$, the $X_{\textnormal{per}}$ denotes the space of periodic functions that locally belong to $X$, for example: $W^{1,p}_{\textnormal{per}}(\square):=\{f\in W^{1,p}_{\textnormal{loc}}(\mathbb{R}^{d})\,:\,\text{$f$ is periodic with periodicity cell $\square$}\}$.
@@ 2 Wave equation
We emphasise that $\Gamma_{\varepsilon}$ and $\mathcal{G}$ are transformations with respect to the spatial variables, whereas $\operatorname{\mathcal{L}}_{\nu}$ acts on the temporal variable and $\operatorname{\mathcal{L}}_{\nu}^{-1}$ on the frequency variable $\lambda=\nu+k$, $k\in\mathbb{R}$. Let us determine
@@
Fix $\varepsilon>0$ and $f\in H^{2}_{\nu}(\mathbb{R};L^{2}(\mathbb{R}^{d}))$, and suppose (2.3) holds. Let $u_{\varepsilon}$ solve (2.1), $U^{\varepsilon}=(\operatorname{\partial_{t}}u_{\varepsilon},a(\tfrac{x}{\varepsilon})\nabla u_{\varepsilon})^{\bot}$ solve (2.2), $u_{0}$ solve (2.16) and $U^{0}=(\operatorname{\partial_{t}}u_{0},a_{0}\nabla u_{0})^{\bot}$ solve (2.18). Then, the following inequalities hold:
@@
By noting that $U^{(1)}-V\in E_{\theta}$ and $A^{(1)}_{\theta}$ is skew-adjoint, one arrives at
@@
Furthermore more, by the properties of the Fourier transform
@@
$\mathcal{F}(\gamma(\tfrac{\cdot}{\varepsilon})\psi))=\langle\gamma\rangle\mathcal{F}\psi+\mathcal{F}\left(\psi\sum_{n\neq 0}\gamma_{n}e^{\mathrm{i}2\pi n\cdot\frac{\cdot}{\varepsilon}}\right).$
@@ 3 Heat equation
Similarly to the case of the wave equation, we employ that $\operatorname{\partial_{t}}$ is boundedly invertible in $L^{2}_{\nu}(\mathbb{R}^{d};L^{2}(\mathbb{R}^{d}))$ for $\nu>0$, (3.7), that $b$ is uniformly positivity with (3.8) as well as (3.7) and (3.9) to deduce the following set of estimates.
@@
where $M=\left(\begin{matrix}\lambda&0\\ 0&b^{-1}\end{matrix}\right)$. Now, from (3.12) and (3.15) it follows that
@@
where $d(t):=\tfrac{t}{|\lambda+t|}$, and $e(t):=\tfrac{t}{|\lambda+t^{2}|}$, for $t\in[0,\infty)$. With $\lambda=\mathrm{i}t+k$ it is straightforward to see that $d(t)=t\big{(}(\nu+t)^{2}+k^{2}\big{)}^{-1/2}\leq 1$ for all $t\in[0,\infty)$. One can readily show that $e(t)=t\big{(}(\nu+t^{2})^{2}+k^{2}\big{)}^{-1/2}\leq e(|\lambda|^{-1/2})\lesssim|\lambda|^{-1/2}$ for all $t\in[0,\infty)$. Putting the above results together gives
@@ 4 Thermo-elastic system of equations
here, $\varepsilon>0$, $u_{\varepsilon},v_{\varepsilon}$ are the unknown solutions to be determined, $f$, $g$ are given, the coefficients $a$, $b$, are matrix-value periodic functions and $\gamma$ is a scalar-valued periodic function; all with fundamental periodicity cell $\square:=[-\tfrac{1}{2},\tfrac{1}{2})^{d}$. We additionally assume $x\mapsto\gamma(\tfrac{x}{\varepsilon})$ (and therefore $x\mapsto\overline{\gamma}(\tfrac{x}{\varepsilon})$) is a multiplier in $H^{1}$, i.e., there is $C\geq 0$ such that
@@
This informs us that $U^{\varepsilon}-U^{0}$ solves a the homogenised evolutionary system with $\varepsilon$-small right-hand-side. Indeed, Theorem 4.1 implies

### v2

@@ 1 Introduction
This article is concerned with the quantitative homogenisation of evolutionary equations (in the sense of Picard) with rapidly oscillating periodic spatial coefficients
@@
where $M$ and $N$ are bounded and positive coefficient matrices, the small parameter $\varepsilon$ is the spatial period and $A$ is some first order differential operator. The usual, now classical, ‘qualitative’ homogenisation theory states that the solution $U_{\varepsilon}$ converges (in some function space), as $\varepsilon$ tends to zero, to $U_{0}$ the solution to an evolutionary equation with constant coefficients
@@
for $U^{0}$ the solution to the homogenised evolutionary heat system
@@
$\partial_{t}u_{0}-\mathrm{div}(b_{0}\nabla u_{0})=f,\quad t\in\mathbb{R},\,x\in\mathbb{R}^{d}.$
As in Section 2, we additionally have the corrector estimate
@@
Upon comparing these estimates with the wave equation error estimates, we see an improvement of the upper bound in the sense that one has a rougher norm for $f$ with respect to the time variable. This is due to the existence of maximal regularity estimates for the evolutionary heat equation (cf. Lemma 3.4) which are utilised when quantitatively justifying the fibre homogenisation approach.
@@
For a given Hilbert space $H$ with inner product $\langle\cdot,\cdot\rangle_{H}$ and norm $\|\cdot\|_{H}:=\langle\cdot,\cdot\rangle^{1/2}$, we denote $L(H)$ to be the space of bounded linear operators in $H$,
@@
and $H_{\nu}^{1}(\mathbb{R};H)$ is the space of $H$-valued, weakly differentiable $L_{2,\nu}(\mathbb{R};H)$-functions with distributional derivative representable as an $L^{2}_{\nu}(\mathbb{R};H)$-function. $L^{2}_{\nu}(\mathbb{R};H)$ and $H^{1}_{\nu}(\mathbb{R};H)$ are Hilbert spaces with the standard inner products. For a given function space $X$, $X_{\textnormal{per}}$ denotes the space of periodic functions that locally belong to $X$, for example: $W^{1,p}_{\textnormal{per}}(\square):=\{f\in W^{1,p}_{\textnormal{loc}}(\mathbb{R}^{d})\,:\,\text{$f$ is periodic with periodicity cell $\square$}\}$.
@@ 2 Wave equation
We emphasise that $\Gamma_{\varepsilon}$ and $\mathcal{G}$ are transformations with respect to the spatial variables, whereas $\operatorname{\mathcal{L}}_{\nu}$ acts on the temporal variable and $\operatorname{\mathcal{L}}_{\nu}^{-1}$ on the frequency variable $\lambda=\nu+\mathrm{i}k$, $k\in\mathbb{R}$. Let us determine
@@
Fix $\varepsilon>0$ and $f\in H^{2}_{\nu}(\mathbb{R};L^{2}(\mathbb{R}^{d}))$, and suppose (2.3) holds. Let $u_{\varepsilon}$ solve (2.1), $U^{\varepsilon}=(\operatorname{\partial_{t}}u_{\varepsilon},a(\tfrac{x}{\varepsilon})\nabla u_{\varepsilon})^{\top}$ solve (2.2), $u_{0}$ solve (2.16) and $U^{0}=(\operatorname{\partial_{t}}u_{0},a_{0}\nabla u_{0})^{\top}$ solve (2.18). Then, the following inequalities hold:
@@
By noting that $U^{(1)}-V\in E_{\theta}$ and $A^{(1)}_{\theta}$ is skew-self-adjoint, one arrives at
@@
Furthermore, by the properties of the Fourier transform
@@
$\mathcal{F}(\gamma(\tfrac{\cdot}{\varepsilon})\psi)=\langle\gamma\rangle\mathcal{F}\psi+\mathcal{F}\left(\psi\sum_{n\neq 0}\gamma_{n}e^{\mathrm{i}2\pi n\cdot\frac{\cdot}{\varepsilon}}\right).$
@@ 3 Heat equation
Similarly to the case of the wave equation, we employ that $\operatorname{\partial_{t}}$ is boundedly invertible in $L^{2}_{\nu}(\mathbb{R}^{d};L^{2}(\mathbb{R}^{d}))$ for $\nu>0$, (3.7), that (the real part of) $b$ is uniformly positive with (3.8) as well as (3.7) and (3.9) to deduce the following set of estimates.
@@
where $M=\left(\begin{matrix}\lambda&0\\ 0&b^{-1}\end{matrix}\right)$. Now, from (3.12) and (3.15) it follows that (cf. Remark 2.7)
@@
where $d(t):=\tfrac{t}{|\lambda+t|}$, and $e(t):=\tfrac{t}{|\lambda+t^{2}|}$, for $t\in[0,\infty)$. With $\lambda=\nu+\mathrm{i}k$ it is straightforward to see that $d(t)=t\big{(}(\nu+t)^{2}+k^{2}\big{)}^{-1/2}\leq 1$ for all $t\in[0,\infty)$. One can readily show that $e(t)=t\big{(}(\nu+t^{2})^{2}+k^{2}\big{)}^{-1/2}\leq e(|\lambda|^{-1/2})\lesssim|\lambda|^{-1/2}$ for all $t\in[0,\infty)$. Putting the above results together gives
@@ 4 Thermo-elastic system of equations
here, $\varepsilon>0$, $u_{\varepsilon},v_{\varepsilon}$ are the unknown solutions to be determined, $f$, $g$ are given, the coefficients $a$, $b$, are matrix-valued periodic functions and $\gamma$ is a scalar-valued periodic function; all with fundamental periodicity cell $\square:=[-\tfrac{1}{2},\tfrac{1}{2})^{d}$. We additionally assume $x\mapsto\gamma(\tfrac{x}{\varepsilon})$ (and therefore $x\mapsto\overline{\gamma}(\tfrac{x}{\varepsilon})$) is a multiplier in $H^{1}$, i.e., there is $C\geq 0$ such that
@@
This informs us that $U^{\varepsilon}-U^{0}$ solves the homogenised evolutionary system with $\varepsilon$-small right-hand side. Indeed, Theorem 4.1 implies